OpenStudy (anonymous):

The equation of a line is x+5y=12 What is the y-intercept of the line?

OpenStudy (misssunshinexxoxo):

X+5y=12 5y=12-x y=(12-x)/5 y=12/5-x/5 The intercept is 12/5

OpenStudy (anonymous):

Thank you!!